Laserfiche WebLink
preserved for analysis by the geologist using standard field <br /> protocol _ No odors or evidence of contamination were noted, and <br /> PID meter readings were all zero. Drill cuttings & unused core <br /> materials were stored as noted above The surface was covered <br /> pending grouting operations which would be done after completion <br /> of the work. Refer to PLATE_XI for more complete details <br /> After decontamination of all equipment, the rig was mobilized on <br /> soil_borzng_"E" as shown on PLATE VIII _ Cores were taken at 5 <br /> foot intervals from 6' to 35' 6" using the same protocol described <br /> earlier Again, there were no odors or visible evidence of <br /> contamination, and PID readings of bagged core material were all <br /> zero Refer to PLATE XII for details of lithology, blow counts, <br /> core recovery, and other pertinent data <br /> After decontamination, the rig was mobilized on soil-boring- "A" <br /> as shown on PLATE VIII, and augers were advanced and cores taken <br /> at approximately 5' intervals from 6' below grade to bottom at <br /> 36' 6" Core materials were described, sealed, and preserved <br /> according to standard field protocol; no odors, visible contami- <br /> nation, or PID readings were noted Refer to PLATE XIII for <br /> lithology, blow counts, & other data <br /> • Upon completion of the five soil borings, they were thoroughly <br /> grouted from bottom to the surface with a cement/bentonite slu ^ry <br /> under the direction of the representative of the San Joaquin Co <br /> EHD All spoils and unused core materials were placed in sealed, <br /> marred drums & stored on site pending results of laboratory <br /> analyzes <br /> The preserved soil samples were transported by the geologist, Wm <br /> Hunter, under chain of custody to FGL Environmental laboratory <br /> with instructions to analyze for BTEXEPA 8020) , TPH-G (EPA <br /> 8015M ) Per agreement with M. Infurna off the County EHD, total <br /> lead, (EPA 7420, ) was to be run only on the sample with highest <br /> `i'PH reading. <br /> On October 13, 1994, a Spectrum Exploratio tr <br /> restricted clearance mast was mobilized a oil borkpg_"F" which <br /> is located at the east side of the site of the remove as <br /> shown on PLATE VIII The rig was equipped with sufficient clean <br /> 5' lengths of 8 1/4" OD hollow-stem augers to complete the work <br /> without the necessity of cleaning flights Only one 24" <br /> California Modified split-spoon core barrel was available, it was <br /> thoroughly cleaned with TSP & clean water prior to each use. <br /> Drilling began at 0915 under the direction of Wm_ Hunter, RG <br /> Drill cuttings from the surface to 5' consisted of soft, loose <br /> brown loam, it appeared to be backfill material The clean core <br /> barrel filled with 2" brass tubes was lowered into the augers and <br /> 7 <br />